In a very violent reaction called the thermite reaction, aluminum metal reacts with iron (III) oxide to form iron metal and aluminum oxide.


Fe2O3 + AI → Fe + Al₂O₃


What mass of aluminum will react 150 g of iron (III) oxide?



1
Expert's answer
2021-05-25T07:51:10-0400

Solution:

Balanced chemical equation:

Fe2O+ 2AI → 2Fe + Al2O3


The molar mass of Fe2O3 is 159.69 g/mol.

Therefore,

(150 g Fe2O3) × (1 mol Fe2O3 / 159.69 g Fe2O3) = 0.9393 mol Fe2O3


According to the equation above: n(Fe2O3) = n(Al)/2

n(Al) = 2 × n(Fe2O3) = 2 × 0.9393 mol = 1.8786 mol


The molar mass of Al is 26.98 g/mol.

Therefore,

(1.8786 mol Al) × (26.98 g Al / 1 mol Al) = 50.68 g Al


Answer: 50.68 grams of aluminum (Al) will react.


OR (short form solution)

According to the equation above:

(150 g Fe2O3)×(1 mol Fe2O3/159.69 g Fe2O3)×(2 mol Al/1 mol Fe2O3)×(26.98 g Al/1 mol Al) = 50.68 g Al